# Combinatorial Selection with Costly Information

### v1

@@ Appendix C Omitted Proofs from Section 5
$\alpha\leq\max_{i}\min(1+\lambda_{i},\frac{1}{\lambda_{i}})\leq\max_{x\in(0,1)}\min(1+x,\frac{1}{x})=\phi.$
@@
Appendix D Omitted Proofs from Section 6
In this chapter of the appendix, we present all omitted proofs from Section 6, in particular the proof of Theorem 6.2. We restate all our results for the reader’s convenience.
@@ Appendix D Omitted Proofs from Section 6
See 6.2 For ease of notation, let $k:=\alpha+1$ and $B:=2^{k^{2}}$. We consider the alternative with weighing cost $c=1$ and random cost $X$ that is continuously distributed in interval $[1,B]$ and has two point masses on 00 and $(k+1)B$, namely:
@@
Fix any committing policy $\pi$ and let $\ell(x):=\operatorname{E}\left[X\in I\right]$ where $I\in\mathcal{I}^{\pi}$ is the unique interval of the policy’s partition that contains $x$. Clearly, $\ell(x)$ is a lower bound to $\rho^{\pi}(x)$ and it is also a non-decreasing function. Thus, in order to prove impossibility of pointwise approximation for any $\alpha^{\prime}<\alpha$, it suffices to show that there exists some $x\in\mathcal{X}$ for which
@@
We begin by considering the committing policy $\pi$ that does not perform any weighings. Then, we simply have $\ell(x)=\rho^{\pi}(x)=\mu$ for all $x\in\mathcal{X}$ and since $\rho^{*}(x)\geq g$ for all $x\in X$, and the claim follows from $\alpha\leq\mu/g$. Next, we consider any committing policy that performs weighings, and let $t$ be the maximum threshold that it uses. Clearly, $t<(k+1)B$ otherwise there is no point in the weighing. This means that the final interval in $\mathcal{I}^{\pi}$ will be $I_{\infty}:=(t,\infty)$. Let $\mu_{\infty}:=\mu(I_{\infty})=\operatorname{E}\left[X|X>t\right]$. We distinguish between the following cases:
  1. 1.
@@
Appendix E The Maximization Setting
E.1 Amortization for Markov Chains
The cost amortization of a Markov chain is defined in the same manner, with the difference that instead of increasing the terminal state values to obtain surrogate costs, we now decrease them to obtain surrogate values. Furthermore, the index of a state now corresponds to the maximum surrogate value among its downwards terminal states instead of the minimum. Formally:
Definition 12 (Cost amortization for maximization settings).
A cost amortization for a Markov chain $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$ is a non-negative vector $b=\{b_{st}\}_{s\in S,t\in T(s)}$ with the property that
$\sum_{t\in T(s)}p(t)b_{st}=p(s)c(s)$
for all states $s\in S$. Based on an amortization, we define:
  • The amortized value of a terminal state $t\in T$ as $\rho_{b}(t):=v(t)-\sum_{s:t\in T(s)}b_{st}$.
  • The surrogate value of the Markov chain $\mathcal{M}$ as the random variable $\rho_{\mathcal{M},b}$ that takes on value $\rho_{b}(t)$ for $t\in T$ with probability $p(t)$.
  • The index of a state $s\in S$ of the Markov chain $\mathcal{M}$ as $I_{\mathcal{M},b}(s)=\max_{t\in T(s)}\rho_{b}(t)$.
Lemma E.1 (Markov chain upper upper for maximization settings).
Consider any matroid-max-CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ over Markov chains and let $b_{i}$ be any cost amortization of $\mathcal{M}_{i}$ with corresponding surrogate cost $\rho_{i}:=\rho_{\mathcal{M}_{i},b_{i}}$ for all $i\in[n]$. Then, the expected utility of any algorithm for $\mathbb{I}$ is at most $\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}\rho_{i}\right]$.
Up next, we extend our definition of water filling amortization. Like before, we start from the terminal states, and define their total cost share to be equal to their value $v(t)$ and then proceed over non-terminal states in decreasing order of level. The difference is that each non-terminal state $s\in S$ distributes its total cost $c(s)$ across its downstream terminals $T(s)$, starting from the terminals with the maximum current total value, until the equation $\sum_{t\in T(s)}p(t)b_{st}=p(s)c(s)$ is satisfied. In other words, instead of increasing the cost of the minimum-cost terminal, we now decrease the value of the maximum-value terminal. Thus, a more suitable term that we will be using in the maximization setting will be water draining. We now use $W^{*}_{\mathcal{M}}$ to denote the water draining surrogate value of a Markov chain $\mathcal{M}$, and $I^{*}_{\mathcal{M}}(s)$ to denote the water draining index of a state $s$ in $\mathcal{M}$.
Definition 13 (Water Draining Index policy).
The Water Draining Index policy for a matroid-max-CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ chooses at every step the Markov chain $i^{*}=\operatorname{argmax}_{i\in\mathcal{F}_{S}}I^{*}_{i}(s_{i})$, where $s_{i}$ is the current state of each Markov chain $\mathcal{M}_{i}$; $S$ is the set of terminated (selected) Markov chains; and $\mathcal{F}_{S}=\{i:S\cup\{i\}\in\mathcal{F}\}$.
From the same observations as in the proof of Lemma 3.1, it immediately follows that this algorithm achieves two desired properties: it always selects the feasible set of Markov chains with maximum total surrogate value, and whenever it advances a state $s\in S$ that contributes to the surrogate value of a terminal state $t\in T(s)$ (i.e. $b_{st}>0$) and nature realizes $t$, the algorithm will select it. From this, and the fact that the algorithm that greedily adds the maximum weight feasible element is optimal for the matroid independent set problem, the following counterpart to Theorem 3.2 follows, establishing the optimality of the Water Draining Index policy for matroid-max-CICS.
Theorem E.2.
For any matroid-max-CICS instance $\mathbb{I}=(\mathcal{M}_{1},\cdots,\mathcal{M}_{n},\mathcal{F})$ over Markov chains, the expected utility of the water draining index policy is equal to $\operatorname{E}\left[\max_{S\in\mathcal{F}}\sum_{i\in S}W^{*}_{\mathcal{M}_{i}}\right]$. The policy is therefore optimal for instance $\mathbb{I}$.
E.2 Optimality Curves
The definition of a local game $(\mathcal{M},y)$ naturally extends to the maximization setting; at any step, the decision maker can either advance the Markov chain $\mathcal{M}$ at a cost (until it reaches a terminal state in which case it may collect its value and terminate), or collect the reward $y$ of the outside option and terminate. Like before, we use $f_{\mathcal{M}}(y)$ to denote the utility of the optimal adaptive policy in this game. From Theorem E.2, we immediately have that
Definition 14 (Water draining surrogate values for MDPs).
Let $\mathcal{M}$ be an MDP with optimality curve $f_{\mathcal{M}}$. The Water Draining Surrogate Value for $\mathcal{M}$ is the random variable $W^{*}_{\mathcal{M}}$ generated by picking a value from the CDF $\frac{d}{dy}f_{\mathcal{M}}(y)$. That is, $W^{*}_{\mathcal{M}}$ is the random variable satisfying $f_{\mathcal{M}}(y)=\operatorname{E}\left[\max(y,W^{*}_{\mathcal{M}})\right]$ for all $y\in\real$.
E.3 Amortization for MDPs
Lemma E.3.
For any MDP $\mathcal{M}=(S,\sigma,A,c,\mathcal{D},V,T)$, there exists an amortized cost function $\rho:T\mapsto\Delta(\real)$ mapping terminal states to distributions over costs and a non-negative cost sharing vector $b=\{b_{st}\}_{s\in S,t\in T(s)}$, such that for all committing policies $\pi\in\mathcal{C}\left(\mathcal{M}\right)$, generating a Markov chain $\mathcal{M}^{\pi}$ with states $S_{\pi}\subseteq S$, terminal states $T_{\pi}\subseteq T\cap S_{\pi}$, and a distribution $p_{\pi}$ over them, the following properties hold:
  1. 1.
    Action Independence. The water draining surrogate cost $W^{*}_{\mathcal{M}}$ corresponds to sampling a terminal state $t\sim p_{\pi}$ and then sampling from distribution $\rho(t)$.
  2. 2.
    Cost Sharing. For all $t\in T_{\pi}$, $\operatorname{E}\left[\rho(t)\right]=v(t)-\sum_{s\in S_{\pi}:t\in T_{\pi}(s)}b_{st}$.
  3. 3.
    Cost Dominance. For all $s\in S_{\pi}$, $\sum_{t\in T_{\pi}(s)}p_{\pi}(t)b_{st}\leq p_{\pi}(s)c(a)$ for the unique action $a\in A(s)$ chosen by $\pi$.

### v2

@@ Appendix F Omitted Proofs from Section 5 (Partial Inspection PB)
$\alpha\leq\max_{i}\min(1+\lambda_{i},\frac{1}{\lambda_{i}})\leq\max_{x\in(0,1)}\min(1+x,\frac{1}{x})=\varphi.$
F.2 Proof of Lemma 5.3
Recall that we have already defined the opening index $g^{o}$ and peeking index $g^{p}$ of a PI $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$ as the solutions to equations $c^{o}=\operatorname{E}_{X\sim\mathcal{D}}\left[(g^{o}-X)^{+}\right]$ and $c^{p}=\operatorname{E}_{X\sim\mathcal{D}}\left[(g^{p}-X-c^{o})^{+}\right]$ respectively. Observe that if the optimal adaptive policy prefers opening over peeking the box for some outside option $y$, then the same will be true for any $y^{\prime}>y$. Also, for $y=0$ peeking is clearly preferable to opening. Thus, the optimality curves $f^{o}_{\mathcal{B}}(y)$ and $f^{p}_{\mathcal{B}}(y)$ will have a unique intersection (up to an interval). We use $\tau$ to denote this intersection; formally, $\tau$ is the maximal solution to equation
This implies that unless $g^{p}<g^{o}<\tau$, the opening action will dominate the peeking action for all values of the outside option $y$ for which accepting it isn’t optimal; as a consequence, such instances trivially admit a $1$-local approximation by committing to the opening action. From now on, we assume that $g^{p}<g^{o}<\tau$ and prove the two missing claims from Section 5 that complete the proof of Lemma 5.3.
See 1
Proof.
By concavity, we have that $f^{o}_{\mathcal{B}}(y)\leq\alpha f^{o}_{\mathcal{B}}(\frac{y}{\alpha})$ for all $y\in\real$ and $\alpha\geq$ 1; thus, we only need to verify the condition for $y$ such that $f_{\mathcal{B}}(\frac{y}{\alpha})=f^{p}_{\mathcal{B}}(\frac{y}{\alpha})$. In other words, we only need to verify
$\min\{y,f^{o}_{\mathcal{B}}(y)\}\leq\alpha f^{p}_{\mathcal{B}}(\frac{y}{\alpha})\;\;\;\;\forall y\in[\alpha g^{p},\alpha\tau].$
For start, we will show that the condition holds for $\alpha=g^{o}/g^{p}$; notice that for this parameter, $f^{o}_{\mathcal{B}}(y)\leq y$ for all $y\in[\alpha g^{p},\alpha\tau]$ and thus we can now write our condition as
$D_{\alpha}(y):=\alpha f^{p}_{\mathcal{B}}(\frac{y}{\alpha})-f^{o}_{\mathcal{B}}(y)\geq 0\;\;\;\;\forall y\in[\alpha g^{p},\alpha\tau].$
Let $F(\cdot)$ denote the CDF of distribution $\mathcal{D}$. By definition of $f^{o}_{\mathcal{B}}(y)$ and $f^{p}_{\mathcal{B}}(y)$, we have
and by taking the derivative of $D_{\alpha}(\cdot)$, we immediately obtain that $D_{\alpha}(\cdot)$ is a weakly increasing function of $y$ and thus the condition only needs to hold on $y=\alpha g^{p}$. Observe that for $\alpha=g^{o}/g^{p}$ and $y=\alpha g^{p}=g^{o}$, we have
$D_{\alpha}(y)=\frac{g^{o}}{g^{p}}\cdot f^{p}_{\mathcal{B}}(g^{p})-f^{o}_{\mathcal{B}}(g^{o})=g^{o}-g^{o}=0$
and thus we obtain that the opening action always achieves a $g^{o}/g^{p}$ local approximation.
The final step of the proof will be to show that
$\frac{g^{o}}{g^{p}}\leq\frac{c^{o}}{c^{p}}\cdot\big{(}1-\frac{c^{o}}{g^{p}}\big{)}$
from which the claim follows.
See 2
Proof.
By concavity, we have that $f^{p}_{\mathcal{B}}(y)\leq\alpha f^{p}_{\mathcal{B}}(\frac{y}{\alpha})$ for all $y\in\real$ and $\alpha\geq$ 1; thus, we only need to verify the condition for $y$ such that $f_{\mathcal{B}}(\frac{y}{\alpha})=f^{o}_{\mathcal{B}}(\frac{y}{\alpha})$; notice that this corresponds to $y\geq\alpha\tau>g^{p}$ and thus $f^{p}_{\mathcal{B}}(y)\leq y$. In other words, we only need to verify
$f^{p}_{\mathcal{B}}(y)\leq\alpha f^{o}_{\mathcal{B}}(\frac{y}{\alpha})\;\;\;\;\forall y\geq\alpha\tau.$
by concavity of $f^{p}_{\mathcal{B}}(\cdot)$. Next, observe that
$h(\frac{c^{o}}{\alpha-1})\leq\frac{\alpha\cdot c^{o}-c^{p}}{\alpha-1}$
and for $\alpha=1+\frac{c^{p}}{c^{o}}$, this statement is equivalent to
$h(\frac{c^{o}\cdot c^{o}}{c^{p}})\leq\frac{c^{o}\cdot c^{o}}{c^{p}}$
which is clearly true, by definition of $h(\cdot)$.
Appendix G Omitted Proofs from Section 6 (Additive PB)
G.1 Proof of Theorem 6.2
See 6.2 We will separately prove the theorem for $k=2$ and $k\geq 3$.
The $k=2$ case.
We use $(X,c_{x})$ and $(Z,c_{z})$ to denote the $k=2$ components of the additive-MDP and $g_{x},g_{z}$ to denote their respective water filling indices. Notice that there are only two (static) committing policies to consider: $\pi_{XZ}$ and $\pi_{ZX}$ depending on whether $X$ is probed first or not. We use $g_{xz}$ and $g_{zx}$ for their respective indices. Using the same structural properties of surrogate costs for additive-MDPes as in the proof of Theorem 6.1, we have that $\pi_{XZ}$ admits an $\alpha_{xz}$-pointwise approximation for
$\alpha_{xz}=\max_{t=(x,z)}\frac{\rho_{xz}(t)}{\rho_{zx}(t)}=\max_{t=(x,z)}\frac{\max(g_{xz},g_{x}+z,z+x)}{\max(g_{zx},g_{z}+x,z+x)}\leq\max_{t=(x,z)}\max(1,\frac{g_{xz}}{g_{zx}},\frac{g_{x}+z}{\max(g_{zx},g_{z}+x,z+x)})$
and for the last term in the maximum, since $x\geq 0$, we have that
$\alpha_{xz}\leq\max_{t=(x,z)}\max(1,\frac{g_{xz}}{g_{zx}},\frac{g_{x}+z}{\max(g_{zx},z)})\leq\max(\frac{g_{xz}}{g_{zx}},1+\frac{g_{x}}{g_{zx}})$
and likewise we obtain that
$\alpha_{zx}\leq\max(\frac{g_{zx}}{g_{xz}},1+\frac{g_{z}}{g_{xz}}).$
Finally, we will now prove that $\min(\alpha_{zx},\alpha_{zx})\leq\varphi$ always. By re-labeling and re-scaling we can assume without loss that $g_{xz}\leq g_{zx}=1$. Furthermore, we have that $g_{xz}\geq g_{x}+g_{z}$. To see why this is true, consider an amortization $\mathcal{M}^{\pi_{xz}}$ where each (height $1$) $Z$-vertex sends the same cost shares $b_{z}$ as in the water filling amortization of $(Z,c_{z})$ to any terminal $(x,z)$, and the (unique) $X$-vertex sends the same cost shares $b_{x}$ as in the water filling amortization of $(X,c_{x})$ to any terminal $(x,z)$ this amortization covers the opening costs by definition, and the surrogate cost of terminal $(x,z)$ is $x+z+b_{x}+b_{z}=\rho_{x}(x)+\rho_{z}(z)\geq g_{x}+g_{z}$. Since water-filling amortization maximizes the minimum surrogate cost across all amortizations, we indeed have $g_{xz}\geq g_{x}+g_{z}$. Thus, we have $g_{x}+g_{z}\leq g_{xz}\leq g_{zx}=1$. These in turn imply that $\alpha_{xz}\leq 1+g_{x}$ and $\alpha_{zx}\leq\frac{1}{g_{x}+g_{z}}\max(1,g_{x}+2g_{z})$. The proof is then completed by showing
$\alpha:=\min\bigg{(}1+g_{x},\frac{1}{g_{x}+g_{z}}\max(1,g_{x}+2g_{z})\bigg{)}\leq\varphi$
for all $g_{x},g_{z}\geq 0$ with $g_{x}+g_{z}\leq 1$. By fixing $g_{x}$, it is not hard to see that the second term gets maximized at either $g_{z}=0$ or $g_{z}=1-g_{x}$. Thus, we have
$\alpha\leq\max_{x\in[0,1]}\min\bigg{(}x+1,\max(\frac{1}{x},2-x)\bigg{)}=\varphi.$
The $k\geq 3$ case.
Let $\pi\in\mathcal{C}\left(\mathcal{M}\right)$ be any commitment of index $g^{\pi}$ for an additive-MDP $\mathcal{M}$ over $k$-components and let $\alpha=\lfloor\frac{k+1}{2}\rfloor$. We will show that there exists a static committing policy $\pi^{\prime}$ such that $g^{\pi^{\prime}}\leq\alpha\cdot g^{\pi}$. Then, by instantiating $\pi$ as any minimum index policy and observing that all static policies are by definition root dominant (i.e. the root state has maximum index across all states), step 1 in the proof of Theorem 6.1 immediately gives us the proof of Theorem 6.2.
Consider the Markov chain $\mathcal{M}^{\pi}$. Like in the proof of Theorem 6.1, we define the index $g^{\pi}(s)$ of a state $s$ as the water filling index of the sub-chain rooted at $s$ if all previously probed values are realized at 00, and its value $V^{\pi}(s)$ as the sum of all realized values in the trajectory from the root to $s$. Observe that these indices and values uniquely determine the water filling cost shares that each state $s$ sends to its downwards terminal states in $\mathcal{M}^{\pi}$; in particular, we have that
$b_{st}$ $=\max\bigg{(}0,g^{\pi}(s)+V^{\pi}(s)-\max_{s^{\prime}\in P(s,t)}(g^{\pi}(s^{\prime})+V^{\pi}(s^{\prime}))\bigg{)}$
$=\max\bigg{(}0,g^{\pi}(s)-\max_{s^{\prime}\in P(s,t)}(g^{\pi}(s^{\prime})+V^{\pi}(s,s^{\prime}))\bigg{)}$
where the inner maximum is taken over all states in the (unique) path from $s$ to $t$, excluding state $s$, and $V^{\pi}(s,s^{\prime})$ denotes the sum of all realizations in the trajectory from $s$ to $s^{\prime}$. The above equality is a direct consequence of the fact that the minimum surrogate cost in the sub-tree of state $s$ immediately after $s$ gets amortized is by definition $g^{\pi}(s)+V^{\pi}(s)$.
We will now define the static policy $\pi^{\prime}$. Observe that any trajectory from the root of $\mathcal{M}^{\pi}$ to a terminal state defines a fixed ordering of the $k$-components; we define $\pi^{\prime}$ as the ordering corresponding to a trajectory of lexicographically maximum index. In other words, $\pi^{\prime}$ probes the same first box as $\pi$, then the second box corresponding to a child of the root with maximum index etc. By re-labeling, we assume that $\pi^{\prime}$ is the static ordering $k\mapsto(k-1)\mapsto\cdots\mapsto 1$.
We now turn our attention to the Markov chain $\mathcal{M}^{\pi^{\prime}}$. We use $h(s)$ to denote the height of state $s$; this is 00 for terminal states and $k$ for the root. Notice that in $\mathcal{M}^{\pi^{\prime}}$, all the states $s$ with $h(s)=j\in[k]$ correspond to the same (remaining) fixed probing order $j\mapsto(j-1)\mapsto\cdots\mapsto 1$, and will have the same index $g^{\pi^{\prime}}(s)=g(j)$, with $g^{\pi^{\prime}}=g(k)$. Furthermore, each state $s$ of $\mathcal{M}^{\pi^{\prime}}$ at height $j\in[k]$ can be mapped to the unique height $j$ state of $\mathcal{M}^{\pi}$ that was chosen by our process; we use $s_{j}$ to denote this state.
We will prove that for all $j\in[k]$, it holds that $g(j)\leq\alpha(j)\cdot g^{\pi}(s_{j})$ for $\alpha(j):=\lfloor\frac{j+1}{2}\rfloor$. This would directly imply that $g^{\pi}=g(k)\leq\alpha(k)\cdot g^{\pi}(s_{k})=\alpha(k)\cdot g^{\pi}$, completing the proof. In order to upper bound the indices $g(j)$ of $\mathcal{M}^{\pi^{\prime}}$, it suffices to argue that the suggested upper bounds generate enough cost shares to cover the amortization of the costs. Since each state $s$ of $\mathcal{M}^{\pi^{\prime}}$ at height $j\in[k]$ shares the same distribution over terminal states with the state $s_{j}$ of $\mathcal{M}^{\pi}$, it will suffice to point-wise compare these cost shares. In other words, it suffices to prove that
$\max\bigg{(}0,\alpha(j)\cdot g^{\pi}(s_{j})-\max_{s^{\prime}\in P^{\prime}(s,t)}(\alpha(h(s^{\prime}))\cdot g^{\pi}(s_{h(s^{\prime})})+V^{\pi^{\prime}}(s,s^{\prime}))\bigg{)}\geq\max\bigg{(}0,g^{\pi}(s_{j})-\max_{s^{\prime}\in P(s_{j},t)}(g^{\pi}(s^{\prime})+V^{\pi}(s_{j},s^{\prime}))\bigg{)}$
for all $j\in[k]$. Our proof will be completed by arguing that our specific definition of $\alpha(j)$ satisfies
$\max_{s^{\prime}\in P^{\prime}(s,t)}\bigg{(}\alpha(h(s^{\prime}))\cdot g^{\pi}(s_{h(s^{\prime})})+V^{\pi^{\prime}}(s,s^{\prime})\bigg{)}\leq\alpha(j)\cdot\max_{s^{\prime}\in P(s_{j},t)}\bigg{(}g^{\pi}(s^{\prime})+V^{\pi}(s_{j},s^{\prime})\bigg{)}$
for all $j\in[k]$, all states $s$ with $h(s)=j$ and all terminal states $t$.
The maximum in the left hand side will get realized at some $s^{\prime}\in P^{\prime}(s,t)$ with $h(s^{\prime}):=j^{\prime}\leq j-1$ (as $h(s)=j$). If $j^{\prime}<j-1$, then we can upper bound the left hand side by $a(j^{\prime})\cdot g^{\pi}(s_{j^{\prime}})+V^{\pi^{\prime}}(s,t)$. Since $\pi^{\prime}$ is a static policy, it is not hard to see that the indices in any trajectory from the root to a terminal will form a decreasing sequence, so $g^{\pi}(s_{j^{\prime}})\leq g^{\pi}(s_{j})$. Furthermore, we have that $V^{\pi^{\prime}}(s,t)=V^{\pi}(s_{j},t)$ and thus our inequality holds as long as $\alpha(j)\geq\alpha(j^{\prime})+1$ which is always the case for $j^{\prime}<j-1$.
It remains to argue about $j^{\prime}=j-1$. In that case, the left hand side is $\alpha(j-1)\cdot g^{\pi}(s_{j-1})+V_{j}$ for some realization $V_{j}$ of the $j$-th component. We will upper bound this through the first term in the right hand side maximum, corresponding to some child vertex $\hat{s}$ of $s_{j}$; in that case, we need to show that
$\alpha(j-1)\cdot g^{\pi}(s_{j-1})+V_{j}\leq\alpha(j)\cdot g^{\pi}(\hat{s})+V_{j}$
and the proof is completed by the fact that our policy selected a trajectory of lexicographically maximum indices and thus $g^{\pi}(s_{j-1})\leq g^{\pi}(s^{\prime})$ as $s_{j-1}$ and $s^{\prime}$ are both children of the same state $s_{j}$ in $\mathcal{M}^{\pi}$.
@@ Appendix G Omitted Proofs from Section 6 (Additive PB)
Appendix H Omitted Proofs from Section 7 (Weighing Scale Problem)
H.1 Proof of Lemma 7.3
See 7.3 We fix a WS alternative $\mathcal{M}$ corresponding to random value $X$ of support $\mathcal{X}:=\mathrm{support}(X)$ and weighing cost $c$. We use $\mu,M,g,h$ and $\kappa$ to denote the parameters of the alternative, as previously defined. In order to develop our pointwise approximation guarantees, we first need to obtain expressions for the surrogate costs.
Surrogate cost of $\mathcal{M}$.
Fact H.1.
$\rho^{*}(x):=\min(h,\max(g,x)).$
Committing policies.
Now, consider any commitment $\pi\in\mathcal{C}\left(\mathcal{M}\right)$. Observe that $\pi$ corresponds to a protocol that determines in advance a decision tree (or a distribution over decision trees) over pre-specified thresholds against which it will weigh $X$, resulting in a Markov chain $\mathcal{M}^{\pi}$. Importantly, any commitment will end up partitioning the support $\mathcal{X}$ of distribution $\mathcal{D}$ into a set of intervals, corresponding to its terminal states. We use $\mathcal{I}^{\pi}$ to denote this set of intervals, and $t(I)$ to denote the terminal state of $\mathcal{M}^{\pi}$ corresponding to interval $I\in\mathcal{I}^{\pi}$. Furthermore, the probability of running the Markov chain $\mathcal{M}^{\pi}$ and ending up in a terminal state $t(I)$ will be precisely $\operatorname{Pr}\left[X\in I\right]$. This allows us to obtain the following characterization of surrogate costs for committing policies.
Fact H.2.
$\rho^{\pi}(x):=W^{*}_{\mathcal{M}^{\pi}}(t(I))$
for the unique interval $I\in\mathcal{I}^{\pi}$ that contains $x$.
$p(s)\cdot c(s)=\sum_{t\in T(S)}p(t)\cdot b_{st}$
$k:=\log\frac{t_{2}}{t_{f}}\leq\log\frac{M}{g}\leq\log\frac{2\mu}{g}$
$W^{*}_{\mathcal{M}^{\pi}}(t(I))\leq\mu(I)+k\cdot g$
To summarize, we have shown that for all $x\in\mathcal{X}$:
$\rho^{\pi}(x)\leq u(x):=kg+2\cdot\min(2\mu,\max(x,g)).$
Notice that the upper bound $u(x)$ is non-decreasing. Now, recall that $\rho^{*}(x)=\min(h,\max(g,x))$ is also a non-decreasing mapping. This, policy $\pi$ will $\alpha$-pointwise approximate $\mathcal{M}$ for
$a=\max_{x\in X}\frac{u(x)}{\rho^{*}(x)}$
H.2 Proof of Theorem 7.2
@@ Appendix H Omitted Proofs from Section 7 (Weighing Scale Problem)
See 7.2 For ease of notation, let $k:=\alpha+1$ and $B:=2^{k^{2}}$. We consider the alternative with weighing cost $c=1$ and random cost $X$ that is continuously distributed in interval $[1,B]$ and has two point masses on 00 and $(k+1)B$, namely:
@@
Fix any commitment $\pi$ and let $\ell(x):=\operatorname{E}\left[X\in I\right]$ where $I\in\mathcal{I}^{\pi}$ is the unique interval of the policy’s partition that contains $x$. Clearly, $\ell(x)$ is a lower bound to $\rho^{\pi}(x)$ and it is also a non-decreasing function. Thus, in order to prove impossibility of pointwise approximation for any $\alpha^{\prime}<\alpha$, it suffices to show that there exists some $x\in\mathcal{X}$ for which
@@
We begin by considering the committing policy $\pi$ that does not perform any weighings. Then, we simply have $\ell(x)=\rho^{\pi}(x)=\mu$ for all $x\in\mathcal{X}$ and since $\rho^{*}(x)\geq g$ for all $x\in X$, and the claim follows from $\alpha\leq\mu/g$. Next, consider any commitment that performs weighings, and let $t$ be the maximum threshold that it uses. Clearly, $t<(k+1)B$ otherwise there is no point in the weighing. This means that the final interval in $\mathcal{I}^{\pi}$ will be $I_{\infty}:=(t,\infty)$. Let $\mu_{\infty}:=\mu(I_{\infty})=\operatorname{E}\left[X|X>t\right]$. We distinguish between the following cases:
@@
Appendix I Omitted Proofs from Section 8 (Optional Inspection PB)
I.1 Proof of Theorem 8.2
See 8.2